OpenStudy (anonymous):

5y^8-125 factor completely

OpenStudy (amistre64):

i smell a diff of cube

OpenStudy (anonymous):

= 5(y^8 - 25) = 5[ (y^4)2 -5^2] resolve further to get all the factors

OpenStudy (amistre64):

it was disguised as a diff of squares tho... im suing my noseologist

OpenStudy (anonymous):

5(y^8-25)=5(y^4+5)(y^2+sqrt(5))(y+\[5^{1/4}\])(y-\[5^{1/4}\])

OpenStudy (anonymous):

5(y^8-25) 5(y^4-5)(y^4+5) by difference 2 squares

OpenStudy (anonymous):

lol i have no idea where to even start , math is so not my thing

OpenStudy (anonymous):

:D ur funny amistre :D

OpenStudy (anonymous):

final ans 5(y^4+5)(y^2+root5)(y^2 - root5) hope i hv resolved all properly....

OpenStudy (anonymous):

Anyone??????

OpenStudy (anonymous):

depend of what you want to factor it over. reals, integers, complex?

OpenStudy (anonymous):

if you are factoring over integers you get \[5(y^8-25)=5(y^4+5)(y^4+5)\]

OpenStudy (anonymous):

typo \[5(y^4+5)(y^4-5)\]

OpenStudy (anonymous):

if reals you get \[5(y^4+5)(y^2+\sqrt{5})(y^2-\sqrt{5})\] \[=5(y^4+5)(y^2+\sqrt{5})(y+\sqrt[4]{5})(y-\sqrt[4]{5})\]

OpenStudy (anonymous):

my guess is this was supposed to be factored over integers, so first answer is right.

OpenStudy (anonymous):

if complex you get a whole nother can of worms